OpenStudy (anonymous):

If r is directly proportional to t and if r = 3/4 when t = 5/8, what is the value of t when r = 9/25? (Answer in decimal form

OpenStudy (anonymous):

somebody please help!

hero (hero):

What happened to the other question?

OpenStudy (anonymous):

i got help lol

OpenStudy (anonymous):

and this ones harder to

hero (hero):

1. Anytime, one variable is directly proportional to another variable, there is always a constant, c, that needs to be found. This relationship can be expressed as an equation. 2. In this case, that equation is r = ct 3. Input the values r = 9/25 and t = 5/8 into the given equation above. Let me know what you get afterwards.

OpenStudy (anonymous):

like the problem im having is with the decimals because i cant do them like i get how to kinda do the equations but not when there using decimals

hero (hero):

Forget about the decimals for now. Just do exactly as I have suggested above.

OpenStudy (anonymous):

so like 9/25 divided by 5/8 would equal c?

hero (hero):

Yes

hero (hero):

And you already know how to find c because you already know how to divide fractions, right?

OpenStudy (anonymous):

haha not when they look like that

hero (hero):

You mean like this: \[\frac{9}{25} \div \frac{5}{8}\]

OpenStudy (anonymous):

should i cross multiply?

hero (hero):

You can only cross multiply if there's an equal sign separating them.

hero (hero):

I suppose you forgot about the reciprocal rule.

OpenStudy (anonymous):

yess i don't know what im doing and i don't know if i ever learned that.

OpenStudy (anonymous):

how do you know so much?

hero (hero):

You mean you never learned that \[\frac{a}{b} \div \frac{c}{d} \rightarrow \frac{a}{b} \times \frac{d}{c}\]

hero (hero):

Trust me, you have no idea how much I know.

OpenStudy (anonymous):

no! ha i just got done the 10th grade last year. they give me really hard stuff because they think i can do it when inreality i cant

OpenStudy (anonymous):

are you like a teacher? or college student?

hero (hero):

Do I have to be a teacher or college student to know how to divide fractions?

hero (hero):

I know eight year olds who can do the same thing.

OpenStudy (anonymous):

can you help me do this study island stuff because i cant. and no

hero (hero):

All you have to do is post your questions and someone will likely help.

OpenStudy (anonymous):

okay